# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v2

@@ Appendix B Proof of Theorem 1
where $n=\#P$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
Finally, we use assumption 2 to align the answer passrate and the accuracy metric. We can adopt the law of total expectation:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=$ $\mathbb{P}_{p}(\mathrm{correct})\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]+\mathbb{P}_{p}(\mathrm{incorrect})\mathbb{E}[\mathrm{Acc|incorrect}]$ (14)
Note that $\mathbb{P}_{p}(\mathrm{correct})=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$, $\mathbb{E}[\mathrm{Acc}|\mathrm{correct}]=1$ and $\mathbb{P}_{p}(\mathrm{incorrect})=1-\mathbb{P}_{p}(\mathrm{correct})$. We also define $g$ as the random guess accuracy performance, thus we have $\mathbb{E}[\mathrm{Acc|incorrect}]=g$. Take these results into Eq. 14, and we have:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]$ $=\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]+(1-\mathbb{E}_{p}[p_{\mathrm{ans}}(C)])*g$ (15)
$=g+(1-g)\mathbb{E}_{p}[p_{\mathrm{ans}}(C)]$ (16)
$=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{2\mu}\right)+o(\mu)$ (17)
theorem A1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
$f(C)=g+(1-g)\exp{(-aC^{-b}-c)}.$ (18)
Appendix C Experiment Settings
@@
Model training. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters. All models are trained on a constant learning rate scheduler with a few-step warmup stage and data with the same distribution. We list model configurations in Table A1.
@@
Evaluation settings. We conducted performance scaling estimation experiments across 8 major LLM evaluation sets. These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities. All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics. Detailed information is provided in Table A2.
@@
Appendix D Performance Prediction for Continue-Pretrained LLMs
@@
The relationship between model parameter scale and the volume of CPT tokens is listed in Table A1.
@@
Results listed in Table A3 show that the proposed COD method achieves an average prediction error of 2.33%.
@@
Appendix E Additional Ablation Studies
E.1 Comparison of Clustering Methods on Extra Evaluation Sets.
We provide additional clustering evaluation results across the remaining evaluation sets in Table A4 and Table A5, which maintain consistency with the conclusions presented in the main text.
Table A4: Clustering performance on advanced task benchmarks (IAD: Intra-cluster Average Distance, OR: Outlier Rate)
Method TriviaQA AGIEval DROP MBPP
[…]
K-Means 0.4388 - 0.4572 - 0.5554 - 0.3383 -
DBScan 0.7039 6.38 0.5591 3.67 0.6651 11.08 0.5060 12.80
MeanShift 0.2521 6.77 0.2886 2.99 0.2507 11.81 0.2167 15.60
Improved-KMeans 0.1239 11.97 0.1536 7.60 0.1428 21.42 0.1667 19.40
Improved-MeanShift 0.1871 11.54 0.2100 11.50 0.1974 19.88 0.1745 21.60
Table A5: Prediction errors on advanced task benchmarks.
Method Mean Max TriviaQA AGIEval DROP MBPP
EE FE EE FE EE FE EE FE EE FE EE FE
K-Means 2.44 3.36 2.97 8.99 2.97 2.46 2.61 2.68 1.66 1.64 2.53 2.67
DBScan 3.04 2.79 6.43 4.36 1.11 0.81 6.43 6.27 3.03 2.66 1.57 1.41
MeanShift 3.21 2.34 4.18 4.90 3.64 4.90 2.63 3.23 4.18 4.00 2.40 1.22
Improved-KMeans 3.38 3.80 5.96 5.56 1.18 1.12 5.96 5.56 3.99 5.24 2.39 3.25
Improved-MeanShift 1.13 1.61 1.58 2.38 1.58 1.64 1.11 2.38 0.26 0.23 1.56 2.19
E.2 Interpolation Method
The comparative results across different benchmarks are shown in Table A6. We report the prediction error between the real performance of a large model and the mapping result.
Prediction Error BBH Math MMLU-pro
Cubic Spline 0.68 1.31 1.37
Cubic Polynomial 3.38 1.12 1.35
Quintic Polynomial 0.18 1.42 1.36
Quartic Polynomial 1.77 1.28 1.35
Furthermore, in Fig. A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
Appendix F Difficulty Distribution of Predictable Subset
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in Fig. A2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subsets, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the MATH evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH exhibits consistent proportions of its predictable subset across varying difficulty levels, as some questions display oscillatory patterns with limited improvement, even with increased computational resources.
@@ Appendix F Difficulty Distribution of Predictable Subset
Appendix G Limitations
@@
Chain-of-thought performance prediction. theorem A1 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [32, 4] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.

### v3

@@ Appendix B Proof of Theorem 1
where $n$ in the number of tasks in $\mathcal{P}$, and $\mu$, $\sigma^{2}$ follow definitions in the proposition.
Finally, we use assumption 3 to align the answer passrate and the accuracy metric.
$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]$ $=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})]$ (20)
$=g+\frac{(1-g)}{n}\sum_{(q,a_{\text{true}})\in P}\exp(-L(q,a_{\text{true}};C))$ (21)
$=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (22)
Rationality of Assumption 3
Assumption 3 is designed to accommodate tasks with finite answer sets. For such tasks, when calculating $\text{Acc}(C)$, possibilities outside the answer set are disregarded. When $p(a_{\text{true}}\mid q,M_{C})$ approaches 0, $\text{Acc}(C)$ is at the level of a random guess, $g$. When $p(a_{\text{true}}\mid q,M_{C})$ approaches 1, $\text{Acc}(C)$ is close to $p(a_{\text{true}}\mid q,M_{C})$. This assumption implies a linear relationship between $\text{Acc}(C)$ and the probability in the $(0,1)$ interval. The theorem itself is also effective for tasks with open answer sets, where the probability of a correct random guess can be assumed to be 0 (i.e., $g=0$).
Appendix C Additional Ablation Studies
C.1 The Criteria for Extrapolatable Subsets
The criteria for fitting the extrapolation formula (Eq.˜4) are designed to ensure the following:
  1. 1
    $a>0$ and $b>0$: These ensure that accuracy is an increasing function of compute. Larger values of a and b signify that task performance scales more distinctly with compute, leading to fitting curves with better scaling properties and differentiability.
  2. 2
    $c>=0$: This ensures the extrapolated curve’s maximum value is less than or equal to $1$. An excessively large c implies that the fitting curve has a very low ceiling, which is characteristic of task subsets with poor scaling properties. These are thus considered non-extrapolatable.
We conducted an ablation study on these parameters, as shown in Tab.˜A1. Starting from our baseline criteria ($a>1$, $b>0.1$, $0\leq c<1$), we individually relaxed the constraints on a, b, and c, and also observed the effect of removing the thresholds entirely.
[…]
Metric / Task Set Baseline ($a>1$, $b>0.1$, $0\leq c<1$) Ablate a ($a>0.5$) Ablate b ($b>0.05$) Ablate c ($0\leq c<0.5$) w/o threshold
[…]
Mean 1.18 1.63 1.18 1.67 1.19 1.63 1.44 1.97 2.39 2.36
Max 2.20 2.23 2.20 2.23 2.20 2.23 2.88 2.99 6.24 5.97
GSM8k 2.20 2.23 2.20 2.23 2.20 2.23 1.82 2.01 2.20 2.23
MATH 1.14 1.28 1.14 1.28 1.14 1.28 1.14 1.28 0.80 0.79
BBH 0.29 1.77 0.29 1.77 0.29 1.77 0.21 2.99 6.24 5.97
TriviaQA 1.58 1.64 1.59 1.93 1.58 1.64 1.07 1.17 2.92 3.52
MBPP 1.56 2.19 1.56 2.19 1.56 2.19 1.56 2.19 1.53 1.11
AGIEval 1.11 2.38 1.11 2.38 1.11 2.38 0.08 1.59 2.22 2.65
DROP 0.26 0.23 0.26 0.23 0.32 0.18 2.72 1.68 2.20 1.57
MMLU-pro 1.27 1.35 1.27 1.35 1.27 1.35 2.88 2.86 1.00 1.03
When the thresholds are removed entirely, the prediction performance degrades significantly. This is because numerous task clusters with poor scaling properties are included in the extrapolation, impairing the overall result. In contrast, individually relaxing the thresholds for a, b, or c still largely preserves the integrity of the filtering criteria. The performance shows only a slight drop compared to the baseline, indicating that while the filtering step is important, our method is not overly sensitive to the specific threshold values.
C.2 Interpolation Method
The comparative results across different benchmarks are shown in Tab.˜A2. We report the prediction error between the real performance of a large model and the mapping result.
Prediction Error BBH Math MMLU-pro
Cubic Spline 0.68 1.31 1.37
Cubic Polynomial 3.38 1.12 1.35
Quintic Polynomial 0.18 1.42 1.36
Quartic Polynomial 1.77 1.28 1.35
Furthermore, in figure˜A1, we plot the mapping process using different interpolation formulas, where the x-axis represents the predictable subset indices and the y-axis represents the full set indices. The red points are the numerical values to be fitted, green points represent predicted values, purple points represent anchor points, and blue points show the actual performance.
When the number of small models used for fitting varies, the polynomial form also needs to be adaptively adjusted to ensure the robustness of the mapping. For instance, in the experiment of predicting the 12B model, the scaling function was adjusted from a quartic polynomial to a cubic polynomial to mitigate overfitting.
C.3 Different Anchor Points in Interpolation Mapping
We test three configurations:
Table A3: Influence of anchor points in the mapping stage on prediction error ($p$). Lower is better.
Method Overall Individual Task Sets
[…]
w/o. anchor 3.96 8.80 2.17 5.46 5.08 1.68 8.80 2.38 4.44 1.68
w. OOD anchor 1.59 2.56 2.22 0.94 1.84 1.04 2.56 1.86 1.18 1.04
w. ID anchor 1.63 2.38 2.23 1.28 1.77 1.64 2.19 2.38 0.23 1.35
Tab.˜A3 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy. This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models. Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models. Additional ablations on the interpolation method (Sec.˜C.2) confirms quartic functions are suitable.
Appendix D Experimental Settings and Training Recipe
@@
Training recipe. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts. All our models are trained from scratch on a corpus of text data. We do not fix the data budget for all models; instead, we maintain a consistent Data-to-CPT (Compute per token) ratio for all models, as mentioned in Sec.˜5.1. We list model configurations in Tab.˜A4.
We adopt the in-house training data that comprises multilingual text corpora, with increased weighting for domains such as STEM, code, and general knowledge, following Llama3 [17], Deepseek-v2 [28], Fineweb-EDU [29], etc. We apply several de-duplication methods and data cleaning mechanisms to each data source to ensure high-quality tokens.
The model architecture is consistent with Llama3.1 [17], incorporating Grouped-Query Attention (GQA) [2], SwiGLU activation function [35], RMSNorm [48] with Pre-normalization, etc. The models are trained using $BF16$ precision with a sequence length of $8192$ and a RoPE [37] base of $500,000$. We employ the AdamW optimizer with $\beta=(0.9,0.95)$, a weight decay of $0.1$, and a dropout rate of $0.1$.
All models are trained on a constant learning rate scheduler with a few-step warmup stage. To determine the learning rate and batch size, we adopt the hyperparameter scaling laws from Liu et al. [28]. Specifically, the optimal learning rate $\eta_{\mathrm{opt}}$, and the optimal batch size $B_{\mathrm{opt}}$ are defined as power laws of the compute, measured in FLOPs: $\eta_{opt}=a_{1}\cdot C^{-b_{1}}$ and $B_{opt}=a_{2}\cdot C^{b_{2}}$, where $a_{1}$, $b_{1}$, $a_{2}$, $b_{2}$ are parameters to be fitted. We perform a grid search on our small models to identify their optimal learning rates and batch sizes, and then extrapolate these findings to the bigger models.
@@
Training Resources. The 7B dense model is trained on 923B tokens, consuming 52,800 H800 GPU-hours. The computational resources used for the other models can be estimated proportionally based on their respective compute requirements.
Evaluation settings and protocol. We conducted performance scaling estimation experiments across eight major LLM evaluation sets. These evaluation sets span a diverse range of capabilities, including Math, Reasoning, Knowledge, Coding, Reading, and general abilities. All pretrained LLMs were evaluated using a few-shot methodology to obtain the performance metrics. Detailed information is provided in Tab.˜A5. Our evaluation methodology aligns with that used for the Llama3 [17] pre-trained models. We assess the models’ capabilities directly through few-shot text completion tasks without any instruction tuning or Supervised Fine-Tuning (SFT). This evaluation method is chosen because even a small amount of SFT data can significantly influence performance on downstream tasks, thereby not reflecting the inherent capabilities of the pre-trained model itself.
Software Framework. All models are trained using the Megatron framework. The evaluation code is an in-house implementation designed to be consistent with the Llama3 [17] evaluation methodology.
@@
Appendix E Additional Prediction Results for a Medium-size Model
[…]
For this experiment, clustering relied exclusively on the passrate@100 data from the 122M to 1.9B source models to prevent data leakage from the target model. The results in Tab.˜A6 show that COD is more robust and less demanding on the size of source models compared to the baselines. For instance, the baseline methods perform poorly on the BBH benchmark, In contrast, COD accurately captures the trend of accuracy growth as compute increases and yields precise extrapolations. We will include a plot of this curve in the main text to visually demonstrate the clear advantage of COD.
The prediction errors of all methods for the 12B model are significantly higher than those for the 70B model reported in Sec.˜5.2. Extrapolation accuracy improves when larger source models are employed for fitting, as a greater number of samples have “emerged”. If the largest source model is excessively small, numerous samples remain “unemerged”, resulting in substantial errors when extrapolating to a significantly larger model.
Table A6: Absolute prediction error (%) on evaluation sets for predicting the performance of the 12B model. Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red). $\downarrow$ indicates lower is better.
Task Sets/Metrics Ground Truth Loss-intermediate End-to-end BNSL COD
Prediction Error Prediction Error Prediction Error Prediction Error
Overall Metrics
Mean - - 14.92 - 10.52 - 17.42 - 4.68
Max - - 23.82 - 24.43 - 28.26 - 10.39
Individual Task Sets
GSM8k 70.47 60.86 9.61 63.66 6.81 80.99 10.52 60.08 10.39
MATH 30.85 41.53 10.68 34.36 3.51 49.26 18.41 28.48 2.37
BBH 60.76 36.94 23.82 36.33 24.43 36.10 24.66 57.751 3.05
TriviaQA 58.84 68.42 9.57 59.88 1.04 87.10 28.26 57.92 0.93
MBPP 47.80 28.53 19.27 36.51 11.29 27.95 19.85 44.50 3.30
AGIEval 44.13 33.91 10.22 35.40 8.74 34.12 10.01 47.53 3.40
DROP 54.35 35.05 19.30 37.19 17.16 36.73 17.62 63.50 9.14
MMLU-pro 37.60 20.74 16.86 26.41 11.20 27.59 10.01 42.42 4.82
Appendix F Performance Prediction for Continue-Pretrained LLMs
@@
The relationship between model parameter scale and the volume of CPT tokens is listed in Tab.˜A4.
@@
Results listed in Tab.˜A7 show that the proposed COD method achieves an average prediction error of 2.33%.
@@
Appendix G Difficulty Distribution of Predictable Subset
We analyze the proportion of predictable subset tasks across different difficulty levels. The difficulty distributions of predictable subset versus complete sets for different evaluation benchmarks are illustrated in figure˜A2. We use the scores from the 12B model as the basis for difficulty classification. The results show that MMLU-pro and GSM8k evaluation sets have larger proportions of predictable subsets, indicating that most questions in these datasets exhibit good performance scaling properties. In contrast, many difficult questions with near-zero scores in the MATH evaluation set fall outside the predictable subset, requiring adjustment during the mapping phase. Meanwhile, BBH exhibits consistent proportions of its predictable subset across varying difficulty levels, as some questions display oscillatory patterns with limited improvement, even with increased computational resources.
@@ Appendix G Difficulty Distribution of Predictable Subset
Appendix H Computational Cost
The extra computational overhead of running COD is not expensive compared to training a series of small models with increasing parameter sizes. The main additional cost comes from performing 100 inference evaluations on the evaluation set for each small model (the cost of clustering algorithm is negligible compared to the inference cost). The computational complexity is $O(TMN)$, where $T$ is the number of evaluation runs, $M$ is the number of tokens for one evaluation, and $N$ is the maximum parameter size of the small models used for prediction. The corresponding token usage is $O(TM)$.
In particular, for an evaluation set requiring $1M$ tokens, a total of $100M$ tokens for small model inference is needed. In our experiments, the training token count for the 12B small model is $1.554T$. Considering that a training token is typically 3 times more costly than an inference token, the additional cost of COD is approximately $\frac{100M}{1.554T*3}\approx 0.002\%$ of the training cost.
Appendix I Limitations
@@
Chain-of-thought performance prediction. theorem˜1 assumes that evaluation sets directly assess models’ ability to provide answers. However, increasingly more evaluations allow models to think before providing answers. Recent works on inference time scaling [36, 5] further demonstrate that for tasks involving mathematics, reasoning, and coding, training models to complete tasks through longer inference computation can significantly improve downstream task performance. In cases where the reasoning process or answers are not unique, the relationship between a model’s answer loss and passrate on a task may not necessarily follow the exponential relationship between the answer loss and the sample passrate. Although our approach maintains its prediction effectiveness in such situations, the theoretical explanation for these cases is insufficient. Therefore, we consider improving prediction methods based on chain-of-thought characteristics and expanding theoretical foundations as future work.